Over the course of the day you do 1.5Mj of mechanical work and lose 28.5Mj of heat while maintaining a constant body temperature.

What mass of food with an energy content of 50 Mj kg -1 would you need to invest in order to avoid weight loss? (2sf and in kg)


1
Expert's answer
2021-05-27T17:48:26-0400

Solution.

"A=1.5MJ;"

"\Delta U=28.5MJ;"

"W_1=50MJ\/kg;"

"\Delta W=A+Q;"

"\Delta W=1.5MJ+28.5MJ=30MJ;"

"m=\dfrac{\Delta W}{W_1};"

"m=\dfrac{30MJ}{50 MJ /kg}=0.6kg;"

Answer: "m=0.6kg."